Find conditions under which a scalar conservation law can have a stationary jump discontinuity.



1
Expert's answer
2021-03-09T03:56:42-0500

A general form of scalar conservation laws with further properties including some basic models and provide examples of computational methods for them. The equations described by


"\u2202_tu+\u2202_xf(u)=0" , t>0, x∈R


in one dimension are known as scalar conservation laws where u=u(t,x) is the conserved quantity


and f=f(u) is the associated flux function depending on t and x.

The jump condition with discontinuous flux function can be given as


"x'(t) = S(u^{x+}, u^{x-})"